# Generative model for financial time series trained with MMD using a signature kernel

### v3

@@ Abstract
Generating synthetic financial time series data that accurately reflects real-world market dynamics holds tremendous potential for various applications, including portfolio optimization, risk management, and large scale machine learning. We present an approach for training generative models for financial time series using the maximum mean discrepancy (MMD) with a signature kernel. Our method leverages the expressive power of the signature transform to capture the complex dependencies and temporal structures inherent in financial data. We employ a moving average model to model the variance of the noise input, enhancing the model’s ability to reproduce stylized facts such as volatility clustering. Through empirical experiments on S&P 500 index data, we demonstrate that our model effectively captures key characteristics of financial time series and outperforms a comparable GAN-based approach. In addition, we explore the application of the synthetic data generated to train a reinforcement learning agent for portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
@@ 1 Introduction
This paper proposes a generative model based on the maximum mean discrepancy (MMD) as defined in [21] with the use of signature kernels proposed by [31] to address these challenges. MMD provides a framework for comparing probability distributions, enabling effective training of the generative model. The signature kernel, derived from the signature transform, offers a powerful tool for capturing nuances of financial time series data. We further enhance the model’s ability to represent stylized facts by employing a moving average model to generate structured noise input, leading to more realistic synthetic data generation.
To demonstrate the effectiveness of our approach, we compare the characteristics of the generated sequence using our MMD approach against two alternative approaches. One approach from [59] uses causal optimal transport (COT) hence named COT-GAN, which is a type of Wasserstein distance that respects causality (see [4]). Another approach from [37], named Conditional Sig-WGAN, is similar to ours in that it also uses a MMD distance but does not first "lift" the sequence to a RKHS (see Section 3.4). We find that although both models are generally able to capture stylised facts of financial time series, there are some improvements in the statistical aspects of the MMD generated sequence. Furthermore, we explore the application of the generated synthetic data for training a reinforcement learning agent on the task of portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
The sections are organised as follows. Section 2 provides an overview of related works in the field of generative models for financial time series that use neural networks. Section 3 details the technical underpinnings behind the methodology. We describe our proposed generative model and the modelling of the noise variance in Section 4. Section 5 presents the experimental results, comparing our approach with the COT-GAN and Conditional Sig-WGAN, and evaluating the performance of the synthetic data generated for portfolio management using reinforcement learning. We also introduce a method to add robustness to the generative model in this section. Finally, Section 6 concludes the paper and discusses potential future research directions.
@@ 2 Related Works
Generative models are a type of statistical model that can be used to generate new data instances. They are especially useful in the financial domain, where they can be used to generate financial time series data. Financial time series data is a sequence of values that are measured at regular intervals, such as daily, weekly, or monthly. Generative models can be used to generate new time series data that is similar to the original data. This is useful for a variety of applications, including portfolio optimization, risk management, and privacy protection. In addition, generative models can be used to backtest trading strategies or serve as training data for machine learning models. Traditional statistical methods for modelling financial time series include autoregressive models, GARCH models, and stochastic volatility models. We focus instead on related works that also use neural networks as the backbone for the generative model.
@@
An even more popular approach is the framework of Generative Adversarial Networks (GANs) pioneered in [20] to train neural networks generators of financial time series. One of the earliest works using this approach is [54] where the authors used multi-layer perceptrons and convolutional networks in the generator and discriminator. The QuantGAN from [55] focuses on the choice of neural network architecture. In particular, an architecture called Temporal Convolutional Networks (TCN) was used to model drift and volatility processes in order to generate samples. [17] further incorporates the attention mechanism with transformers into the network architecture under the GAN framework. One common issue with the GAN framework is that it is known to be exhibit instabilities during training. The authors of QuantGAN report that "training was very irregular and did not converge to a local optimum of the objective function", hence the need for model selection across multiple parameters saved during training.
@@
The closest related work is [27] where the authors also used neural SDEs as the base model and train the generator using a modified version of the MMD with the signature kernel. Similar to our approach, the authors also use the insight from [31] to first lift the path to the canonical feature space of a classical kernel such as the Gaussian kernel. One key difference to the approach in this paper is the distribution of the noise input to the generative model. It is typical in generative models to use independently distributed Gaussian noise for each time step. We use a moving average model to model the variance of the noise which is found to be important in generating autocorrelation in the absolute returns of the generated sequences which is a stylised fact of financial time series. The use of the moving average model is novel to the best of our knowledge.
In terms of applying the generative model to downstream tasks, the authors of [10] simulate limit order book data while in [34], the authors create synthetic data for building or fine tune trading strategies. The strategies are based on data driven machine learning techniques such as neural networks and gradient boosting machines. We will do a similar portfolio management task using reinforcement learning trained on synthetic data created by the generative model. In particular, we look to generate synthetic data of longer time horizons which is another aspect that differentiates our approach from the existing works.
@@ 3.1 Reproducing Kernel Hilbert Space
The primary advantage of using a kernel is that there is a way to compute the result of the inner product without having to explicitly compute the feature map $\phi(x)$. This results in efficient computation despite the fact that the inner product space ${\mathcal{V}}$ may be of high dimension or even infinite dimensional. This is known as the kernel trick (see [52, Section 4]). Every kernel possesses a corresponding unique reproducing kernel Hilbert space (RKHS) (see [44]).
@@
[…]
@@ 3.2 Kernel Two Sample Test
The MMD between two probability measures $P,Q\in{\mathcal{P}({\mathcal{X}})}$ is defined as the supremum of the difference between the expectation of the function evaluated on the two distributions where the supremum is taken over the unit ball of a RKHS $({\mathcal{H}},\langle\cdot\;,\;\cdot\rangle_{\mathcal{H}})$ with respect to the induced norm $\|\cdot\|_{\mathcal{H}}=\sqrt{\langle\cdot\;,\;\cdot\rangle_{\mathcal{H}}}$.
@@
Firstly, by [44, Lemma 3.1], $\mu_{P},\mu_{Q}\in{\mathcal{H}}$. Moreover
@@
$\widehat{\text{MMD}^{2}}(X,Y;k):=\underbrace{\frac{1}{m(m-1)}\sum^{m}_{i=1}\sum^{m}_{j=i,j\neq i}k(x_{i},x_{j})}_{\text{A}}-\underbrace{\frac{2}{m^{2}}\sum^{m}_{i=1}\sum^{m}_{j=1}k(x_{i},y_{j})}_{\text{B}}+\underbrace{\frac{1}{m(m-1)}\sum^{m}_{i=1}\sum^{m}_{j=1,j\neq i}k(y_{i},y_{j})}_{\text{C}}$ (6)
@@ 3.3 Signature of a Path
The kernel two sample tests described in Section 3.2 are typically used for distributions over "static" i.e. non-causal values. However, we are interested in distributions over paths and to capture the time causalities of these path values. We can extend the test to distributions over paths by using the $\widehat{\text{MMD}^{2}}$ defined in equation (6) with $k$ being the signature kernel [9]. Before we define the signature of a path which is a sequence of iterated integrals, we use an abbreviated notation for the integral. For a path ${\mathbf{x}}\in{\mathscr{P}}:=\{{\mathbf{x}}:[0,T]\to{\mathbb{R}}^{d}\}$ where at each time $t\in[0,T]$, ${\mathbf{x}}_{t}:=(x_{t}^{1},\ldots,x_{t}^{d})$ is the value of the path at time $t$, the integral of the path from time 00 to $t$ of asset $i\in\{1,\ldots,d\}$ is denoted as $S({\mathbf{x}})^{i}_{0,t}$.
[…]
@@
where the zeroth term is defined by convention to be 1 and the superscipt runs through all possible combinations of indices $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$ and $k\geq 1$.
@@
    Uniqueness of the signature of a path: ${\mathbf{S}}({\mathbf{x}})_{0,T}$ is unique to ${\mathbf{x}}$ up to a tree-like equivalence. Refer to [22] for the definition of tree-like equivalence. In particular, the signature would be invariant to translations e.g. adding a constant to all the values of the path. Adding a monotone coordinate to the path, such as time, is a sufficient condition for the uniqueness of the signature.
@@
    Invariance to time reparameterisation: If $\tilde{{\mathbf{x}}}$ is a time reparameterisation of ${\mathbf{x}}$ i.e. $\tilde{{\mathbf{x}}}_{t}={\mathbf{x}}_{\psi(t)}$, then ${\mathbf{S}}({\mathbf{x}})_{0,T}={\mathbf{S}}(\tilde{{\mathbf{x}}})_{0,T}$ where $\psi:[0,T]\to[0,T]$ is a strictly increasing, continuous and surjective function. This can be shown with a change of variables in the iterated integrals.
  3. 3.
    Factorial decay of the signature terms [42, Proposition 2.2]: Assuming ${\mathbf{x}}$ is a path of finite total variation. For each $1\leq k\in{\mathbb{N}}$, the norm of the collection of terms of the signature of order $k$ (i.e. $k$ times iterated itegrals) is bounded as follows
@@
  4. 4.
    Universal non-linearity [36, Theorem 2.1] :
    Let $D\subseteq{\mathscr{P}}$ be a compact subset [^3] Refer to [40, Section 8] on the topology of such a set of paths of the space of paths with bounded total variation from $[0,T]$ to ${\mathbb{R}}^{d}$ such that for any ${\mathbf{x}}\in D$, ${\mathbf{x}}_{0}=\mathbf{0}$ and at least one dimension of ${\mathbf{x}}$ is monotone (for uniqueness of the signature). Let $g:D\to{\mathbb{R}}$ be a continuous function. Then, for any $\epsilon>0$, there exists a linear functional $h$ such that for all ${\mathbf{x}}\in D$,
@@
Combining properties 1, 2 and 4 serves also as the motivation for augmenting the path with a time dimension. As the time dimension is non-decreasing, it would ensure that the signature of each time augmented path is unique. Property 2 implies it is only the order of events that matter and not the time scales between the events. Adding a time dimension will therefore allow the signature to capture the time scale related features of the path. The universal non-linearity requires that at least one dimension of the path is monotone which is ensured by the augmentation of the time dimension.
Due to property 3, we can approximate the signature with a truncation at order $m$ and have a bound on the error between the truncated signature and the non-truncated signature.
@@
Indeed, due to property 3, we have for all ${\mathbf{x}}\in{\mathscr{P}},m\in{\mathbb{N}}$ that
@@ 3.4 Signature Kernel
Let $({\mathcal{H}},\langle\cdot\;,\;\cdot\rangle_{\mathcal{H}})$ be the RKHS of a kernel $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$ and ${\mathbf{S}}^{m}$ be the truncated signature map of order $m$. The signature kernel is defined as the inner product ${\mathscr{P}}\times{\mathscr{P}}\ni({\mathbf{x}},{\mathbf{y}})\mapsto{k^{m}_{\text{sig}}}({\mathbf{x}},{\mathbf{y}}):=\langle{\mathbf{S}}^{m}(k_{\mathbf{x}}),{\mathbf{S}}^{m}(k_{\mathbf{y}})\rangle$ of the lifted paths $k_{\mathbf{x}}=(k_{{\mathbf{x}}_{t}})_{t\in[0,T]}$, $k_{\mathbf{y}}=(k_{{\mathbf{x}}_{t}})_{t\in[0,T]}\in{\mathscr{P}}_{\mathcal{H}}:=\{{\mathbf{x}}:[0,T]\to{\mathcal{H}}\}$ i.e. the space of paths evolving in the RKHS ${\mathcal{H}}$ and $k_{z}=k(z,\cdot)$ for $z\in{\mathcal{X}}$.
The kernel that is used to lift the paths is often called a static kernel and explicit computations of ${\mathbf{S}}(k_{\mathbf{x}}),{\mathbf{S}}(k_{\mathbf{y}})\in\prod^{m}_{k=0}{\mathcal{H}}^{\otimes k}$ are avoided with only evaluations of the static kernel $k({\mathbf{x}}_{t},{\mathbf{y}}_{t})$ required. We primarily use the rational quadratic kernel as the static kernel. It is a characteristic kernel and has been reported to perform well in practice (see e.g. [6]).
@@
The rational quadratic kernel can be seen as a weighted sum of an infinite number of Gaussian kernels with different length scales where the squared inverse length scale follows a gamma distribution (see [57, Section 4.2]). The choice of the length scale parameter is important as it determines the smoothness of the kernel function. It is a hyperparameter that needs to be tuned for the specific problem at hand. The larger the length scale, the smoother the kernel function which means it is less sensitive to differences in the inputs. However, if the chosen length scale is too small then even the smallest differences arising from samples from the same distribution will be overemphasised.
The computation of the signature kernel from Definition 7 can be done efficiently using only evaluations of the static kernel on the paths ${\mathbf{x}}$ and ${\mathbf{y}}$ as shown in [31, Algorithm 3]  [^5] There is also a method to compute an untruncated signature kernel by numerically solving a PDE where the solution is equal to the untruncated signature kernel introduced in [47]. However, the MMD failed to converge in our experiments with this method.. The main hyperparameter is the truncation order $m$ of the signature. We modify the algorithm provided in https://github.com/tgcsaba/KSig for our specific use case.

### v4

# Generative modelling of financial time series with structured noise and MMD-based signature learning

@@ Abstract
Generating synthetic financial time series data that accurately reflects real-world market dynamics holds tremendous potential for various applications, including portfolio optimization, risk management, and large scale machine learning. We present an approach that uses structured noise for training generative models for financial time series. The expressive power of the signature transform has been shown to be able to capture the complex dependencies and temporal structures inherent in financial data when used to train generative models in the form of a signature kernel . We employ a moving average model to model the variance of the noise input, enhancing the model’s ability to reproduce stylized facts such as volatility clustering. Through empirical experiments on S&P 500 index data, we demonstrate that our model effectively captures key characteristics of financial time series and outperforms comparable approaches. In addition, we explore the application of the synthetic data generated to train a reinforcement learning agent for portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
@@ 1 Introduction
The signature of a path has been used effectively as a means to train generative models for financial time series using the maximum mean discrepancy (MMD) as defined in [21]. This is done with signature kernels, for example in [27] by using computational methods proposed in [31] and [46]. The prior works in this area typically use identical and independent Gaussian noise as the key stochastic input to these models. This paper proposes enhancing the model’s ability to represent stylized facts by employing a moving average model to generate structured noise input, leading to more realistic synthetic data generation.
To demonstrate the effectiveness of our approach, we compare the characteristics of the generated sequence using our MMD approach against similar approaches using the signature kernel as well as an alternative approach from [58] that uses causal optimal transport (COT), which is a type of Wasserstein distance that respects causality (see [4]). Signature kernel methods have been applied in [36] through the Conditional Sig-WGAN framework, and in [27], where neural SDEs were incorporated into the model architecture. We find that although all the models are generally able to capture most stylised facts of financial time series, there are some improvements in the statistical aspects of our MMD generated sequence. Furthermore, we explore the application of the generated synthetic data for training a reinforcement learning agent on the task of portfolio management, achieving promising results. Finally, we propose a method to add robustness to the generative model by tweaking the noise input so that the generated sequences can be adjusted to different market environments with minimal data.
The sections are organised as follows. Section 2 provides an overview of related works in the field of generative models for financial time series that use neural networks. Section 3 details the technical underpinnings behind the methodology. We describe our proposed generative model and the modelling of the noise variance in Section 4. Section 5 presents the experimental results, comparing our approach with the alternatives, and evaluating the performance of the synthetic data generated for portfolio management using reinforcement learning. We also introduce a method to add robustness to the generative model in this section. Finally, Section 6 concludes the paper and discusses potential future research directions.
@@ 2 Related Works
Generative models are a type of statistical model that can be used to generate new data instances. They are especially useful in the financial domain, where they can be used to generate financial time series data. Financial time series data is a sequence of values that are either measured at regular intervals e.g. index price series, or in some cases at irregular time intervals e.g. limit order books. Generative models can be used to generate new time series data that is similar to the original data. This is useful for a variety of applications, including portfolio optimization, risk management, and privacy protection. In addition, generative models can be used to backtest trading strategies or serve as training data for machine learning models. Traditional statistical methods for modelling financial time series include autoregressive models, GARCH models, and stochastic volatility models. We focus instead on related works that also use neural networks as the backbone for the generative model.
@@
An even more popular approach is the framework of Generative Adversarial Networks (GANs) pioneered in [20] to train neural networks generators of financial time series. One of the earliest works using this approach is [53] where the authors used multi-layer perceptrons and convolutional networks in the generator and discriminator. The QuantGAN from [54] focuses on the choice of neural network architecture. In particular, an architecture called Temporal Convolutional Networks (TCN) was used to model drift and volatility processes in order to generate samples. [17] further incorporates the attention mechanism with transformers into the network architecture under the GAN framework. One common issue with the GAN framework is that it is known to exhibit instabilities during training. The authors of QuantGAN report that "training was very irregular and did not converge to a local optimum of the objective function", hence the need for model selection across multiple parameters saved during training.
@@
The closest related work is [27] where the authors used neural SDEs as the base model and train the generator using a modified version of the MMD with the signature kernel. Similar to our approach, the authors also use the insight from [31] to first lift the path to the canonical feature space of a classical kernel such as the Gaussian kernel. One key difference to the approach in this paper is the distribution of the noise input to the generative model. It is typical in generative models to use independently distributed Gaussian noise for each time step. We use a moving average model to model the variance of the noise which is found to be important in generating autocorrelation in the absolute returns of the generated sequences which is a stylised fact of financial time series. The use of the moving average model is novel to the best of our knowledge. The additional structure in the noise also allows us to quickly adapt the generative model to produce time series that are better aligned to different market environments by only adjusting the noise distribution. This adds efficiency by requiring less data to refit the model and is also quick to implement in practice. The architecture of the neural network is also different to [27] as we use a recurrent neural network (RNN) based architecture instead of a neural SDE. Specifically, we use a Long Short-Term Memory (LSTM) network as the backbone of the generator, which provides a structured way of capturing the temporal dependencies in the data. Finally, we perform direct comparisons between our approach and related works such as the COT-GAN, Conditional Sig-WGAN and the signature kernel trained neural SDEs from [58], [36] and [27] respectively, to demonstrate the effectiveness of our approach.
In terms of applying the generative model to downstream tasks, the authors of [10] simulate limit order book data while in [34], the authors create synthetic data for building or fine tuning trading strategies. The strategies are based on data driven machine learning techniques such as neural networks and gradient boosting machines. We will do a similar portfolio management task using reinforcement learning trained on synthetic data created by the generative model. In particular, we look to generate synthetic data of longer time horizons which is another aspect that differentiates our approach from the existing works.
@@ 3.1 Reproducing Kernel Hilbert Space
The primary advantage of using a kernel is that there is a way to compute the result of the inner product without having to explicitly compute the feature map $\phi(x)$. This results in efficient computation despite the fact that the inner product space $\mathcal{V}$} may be of high dimension or even infinite dimensional. This is known as the kernel trick (see [51, Section 4]). Every kernel possesses a corresponding unique reproducing kernel Hilbert space (RKHS) (see [43]).
@@
Given the set $\mathcal{P}({\mathcal{X}})$} of all probability measures on $\mathcal{X}$} equipped with the Borel $\sigma$-algebra, a kernel $k$ is characteristic if the map [^2] We will use $E_{X\sim P}[\,\cdot\,]$ or $E_{P}[\,\cdot\,]$ to represent the integral $\int_{\mathcal{X}}\,\cdot\;dP(X)$ for brevity. ${\mathcal{P}({\mathcal{X}})}\ni P\mapsto\mu_{P}:=\int_{\mathcal{X}}k(X,\cdot)dP(X)=E_{X\sim P}[k(X,\cdot)]\in{\mathcal{H}}$ is injective. $\mu_{P}$ is called the kernel mean embedding of the probability measure $P$.
@@ 3.2 Kernel Two Sample Test
The MMD between two probability measures $P,Q\in{\mathcal{P}({\mathcal{X}})}$ is defined as the supremum of the difference between the expectation of the function evaluated on the two distributions where the supremum is taken over the unit ball of a RKHS $({\mathcal{H}},\langle\cdot\;,\;\cdot\rangle_{\mathcal{H}})$ with respect to the induced norm $\|\cdot\|_{\mathcal{H}}=\sqrt$$\langle\cdot\;$$\cdot\rangle_{\mathcal{H}}$}.
@@
Firstly, by [RN1023, Lemma 3.1], $\mu_{P},\mu_{Q}\in{\mathcal{H}}$. Moreover
@@
$\widehat{\text{MMD}^{2}}(X,Y;k):=\underbrace{\frac{1}{m(m-1)}\sum^{m}_{i=1}\sum^{m}_{j=1,j\neq i}k(x_{i},x_{j})}_{\text{A}}-\underbrace{\frac{2}{m^{2}}\sum^{m}_{i=1}\sum^{m}_{j=1}k(x_{i},y_{j})}_{\text{B}}+\underbrace{\frac{1}{m(m-1)}\sum^{m}_{i=1}\sum^{m}_{j=1,j\neq i}k(y_{i},y_{j})}_{\text{C}}.$ (6)
@@ 3.3 Signature of a Path
The kernel two sample tests described in Section 3.2 are typically used for distributions over "static" i.e. non-causal values. However, we are interested in distributions over paths and to capture the time causalities of these path values. We can extend the test to distributions over paths by using the $\widehat{\text{MMD}^{2}}$ defined in equation (6) with $k$ being the signature kernel [9]. Before we define the signature of a path which is a sequence of iterated integrals, we use an abbreviated notation for the integral. For a path ${\mathbf{x}}\in{\mathscr{P}}:=\{{\mathbf{x}}:[0,T]\to{\mathbb{R}}^{d}\}$ where at each time $t\in[0,T]$, ${\mathbf{x}}_{t}:=(x_{t}^{1},\ldots,x_{t}^{d})$ is the value of the path at time $t$, the integral of the path from time 0 to $t$ of asset $i\in\{1,\ldots,d\}$ is denoted as $S({\mathbf{x}})^{i}_{0,t}$.
$S({\mathbf{x}})^{i}_{0,t}:=\int_{0<s<T}d{\mathbf{x}}^{i}_{s}={\mathbf{x}}^{i}_{T}-{\mathbf{x}}^{i}_{0}.$ (7)
@@
where the zeroth term is defined by convention to be 1 and the superscript runs through all possible combinations of indices $i_{1},\ldots,i_{k}\in\{1,\ldots,d\}$ and $k\geq 1$.
@@
    Uniqueness of the signature of a path: The signature ${\mathbf{S}}({\mathbf{x}})_{0,T}$ is unique to $\mathbf{x}$} up to a tree-like equivalence. Refer to [22] for the definition of tree-like equivalence. Loosely speaking, there are 3 main instances that are invariant for the signature which correspond to instances where a path integral is invariant. First, a translation, i.e., ${\mathbf{x}}+C$ where $C$ is a constant value, would not change the value of the path integral since $d{\mathbf{x}}=d({\mathbf{x}}+C)$. Second, a time reparameterisation, i.e., ${\mathbf{x}}_{\psi(t)}$ where $\psi:[0,T]\to[0,T]$ is a strictly increasing, continuous and surjective function, would not change the value of the path integral with a simple change of variables. Third, any retracements within the path, i.e., going from point $A$ to $B$ then back to $A$ via the exact same path, would not add any value to the path integral. These are all captured by the definition of a tree-like equivalence.
@@
    Factorial decay of the signature terms [41, Proposition 2.2]: Assuming $\mathbf{x}$} is a path of finite total variation. For each $1\leq k\in{\mathbb{N}}$, the norm of the collection of terms of the signature of order $k$ (i.e. $k$ times iterated integrals) is bounded as follows
@@
  3. 3.
    Universal non-linearity [15, Proposition 3] :
    Let $D\subseteq{\mathscr{P}}$ be a compact subset [^3] Refer to [39, Section 8] on the topology of such a set of paths of the space of paths with bounded total variation from $[0,T]$ to ${\mathbb{R}}^{d}$ such that for any ${\mathbf{x}}\in D$, ${\mathbf{x}}_{0}=\mathbf{0}$ and at least one dimension of $\mathbf{x}$} is a fixed monotone coordinate (for uniqueness of the signature). Let $g:D\to{\mathbb{R}}$ be a continuous function. Then, for any $\epsilon>0$, there exists a linear functional $h$ such that for all ${\mathbf{x}}\in D$,
@@
The above properties also serve as the motivation for augmenting the path with a fixed time dimension, i.e., ${\mathbf{x}}\mapsto(t,{\mathbf{x}})$. A fixed time dimension anchors the path and ensures that the signature of each time augmented path is unique. The universal non-linearity requires that at least one dimension of the path is monotone which is ensured by the augmentation of the time dimension.
Due to property 2, we can approximate the signature with a truncation at order $m$ and have a bound on the error between the truncated signature and the non-truncated signature.
@@
Indeed, due to property 2, we have for all ${\mathbf{x}}\in{\mathscr{P}},m\in{\mathbb{N}}$ that
@@ 3.4 Signature Kernel
Let $({\mathcal{H}},\langle\cdot\;,\;\cdot\rangle_{\mathcal{H}})$ be the RKHS of a kernel $k:{\mathcal{X}}\times{\mathcal{X}}\to{\mathbb{R}}$ and ${\mathbf{S}}^{m}$ be the truncated signature map of order $m$. The signature kernel is defined as the inner product ${\mathscr{P}}\times{\mathscr{P}}\ni({\mathbf{x}},{\mathbf{y}})\mapsto{k^{m}_{\text{sig}}}({\mathbf{x}},{\mathbf{y}}):=\langle{\mathbf{S}}^{m}(k_{\mathbf{x}}),{\mathbf{S}}^{m}(k_{\mathbf{y}})\rangle$ of the lifted paths $k_{\mathbf{x}}=(k_{{\mathbf{x}}_{t}})_{t\in[0,T]}$, $k_{\mathbf{y}}=(k_{{\mathbf{y}}_{t}})_{t\in[0,T]}\in{\mathscr{P}}_{\mathcal{H}}:=\{{\mathbf{x}}:[0,T]\to{\mathcal{H}}\}$ i.e. the space of paths evolving in the RKHS $\mathcal{H}$} and $k_{z}=k(z,\cdot)$ for $z\in{\mathcal{X}}$.
The kernel that is used to lift the paths is often called a static kernel and explicit computations of ${\mathbf{S}}(k_{\mathbf{x}}),{\mathbf{S}}(k_{\mathbf{y}})\in\prod^{m}_{k=0}{\mathcal{H}}^{\otimes k}$ are avoided with only evaluations of the static kernel $k({\mathbf{x}}_{t},{\mathbf{y}}_{t})$ required. This can be done efficiently by approximating the signature kernel with discretised integrals as shown in [31, Algorithm 3] [^5] For an exact computation, there is also [31, Algorithm 6] for higher order computations. Choosing the order to be equal to the truncation level would provide the exact computation. We experimented with the higher order computations which requires significantly more memory but did not see meaningful improvements hence we use the 1st order approximations in our final results. There is also a method to compute an untruncated signature kernel by numerically solving a PDE where the solution is equal to the untruncated signature kernel introduced in [46]. However, the MMD failed to converge in our experiments with this method.. We primarily use the rational quadratic kernel as the static kernel. It is a characteristic kernel and has been reported to perform well in practice (see e.g. [6]).
@@
The rational quadratic kernel can be seen as a weighted sum of an infinite number of Gaussian kernels with different length scales where the squared inverse length scale follows a gamma distribution (see [56, Section 4.2]). The choice of the length scale parameter is important as it determines the smoothness of the kernel function. It is a hyperparameter that needs to be tuned for the specific problem at hand. The larger the length scale, the smoother the kernel function which means it is less sensitive to differences in the inputs. However, if the chosen length scale is too small then even the smallest differences arising from samples from the same distribution will be overemphasised.
The main hyperparameter is the truncation order $m$ of the signature. We modify the algorithm provided in https://github.com/tgcsaba/KSig for our specific use case.